OpenStudy (anonymous):

Can someone help me in this question? @BunnyBree Reiko started a business selling home medical supplies. She spent $5200 to obtain her merchandise, and it costs her $550 per week for general expenses. She earned $900 per week in sales. What is the minimum number of weeks it will take for Reiko to make a profit? Solve the problem.

OpenStudy (samanthagreer):

900w - 550w - 6500 = 0 => 350w = 6500 => w = 18.57 minimum of 19 weeks to make profit

OpenStudy (anonymous):

Thank you! @samanthagreer

OpenStudy (samanthagreer):

no problem can i have a medal please

OpenStudy (anonymous):

no wait that not one of the options :/ @samanthagreer

OpenStudy (anonymous):

that's*

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

where would you start? Ant ideas?

pooja195 (pooja195):

900w - 550w - 6500 = 0 here is what you start off with now you should be able to figure it out :)

pooja195 (pooja195):

@iLorita

OpenStudy (anonymous):

okay... so you subtract like terms right? @pooja195

pooja195 (pooja195):

yes! good job! :D

OpenStudy (anonymous):

so it'll be 350w-6500=0

OpenStudy (anonymous):

then I add 6500 + 0?

pooja195 (pooja195):

well yes and u should have 350w = 6500

pooja195 (pooja195):

and when you divide you will have your answer :D

OpenStudy (anonymous):

I got 18.57

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

15 weeks 14 weeks 13 weeks 12 weeks These are the options @pooja195

OpenStudy (anonymous):

@HannahA

OpenStudy (hannaha):

yes?

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

@HannahA

OpenStudy (hannaha):

I wish I could, but I'm sorry :( I don't really understand this...

OpenStudy (hannaha):

Good luck tho :/

OpenStudy (anonymous):

Do you know anybody that can help? @HannahA

OpenStudy (anonymous):

Thank you :( @HannahA

OpenStudy (hannaha):

You're welcome :) I would recommend some people but most of them are offline

OpenStudy (anonymous):

ok :( @HannahA

OpenStudy (anonymous):

@jskhupmang

OpenStudy (anonymous):

@RANE

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

Can anyone help me?

OpenStudy (butterflydreamer):

i tried doing this question and i can't seem to really get it but i kinda.. used guess and check? (BTW HOW DID PEOPLE GET 6500??? ) So i just wrote Let weeks =w So each week she loses -550w and gains 900w So i just wrote 900 w - 550w > 5200 Then if you sub w= 15 you get 5250 > 5200 so she made a $50 profit... BUT I have no reasoning or explanations for my working out but i tried LOL

OpenStudy (anonymous):

Thank you so much!! lol

OpenStudy (butterflydreamer):

seriously it could be wrong LOL

OpenStudy (anonymous):

Well at least you tried

OpenStudy (butterflydreamer):

true ahahah xD!

OpenStudy (anonymous):

lol cx